Help me solve 4x + x(x-2)=0
Answer by richwmiller(17219)

some website rules you will not like
limit one problem per submission not two not three
no similar problems
limit 4 submissions daily
4x+x*(x-2)=0
deal with the parentheses first
x^2-2x+4x=0
x^2+2x=0
x*(x+2)=0
x=0
x=-2
BTW
Your second problem is incomplete.
There is no equal sign I am not going to guess where it belongs but it might be identical to the first problem.